OpenStudy (aimmebgood):

x^(2) +3xy + 2y^(2) factor. what do you factor?

OpenStudy (anonymous):

you make it x^2+2xy+y^2+xy=(x+y)^2+xy

OpenStudy (aimmebgood):

I still don't understand it.

OpenStudy (anonymous):

3xy=2xy+xy right?

OpenStudy (aimmebgood):

right.

OpenStudy (anonymous):

x^2+2xy+y^2=x^2+xy+xy+y^2=x(x+y)+y(x+y)=(x+y)*(x+y)=(x+y)^2 right?

OpenStudy (aimmebgood):

yeah

OpenStudy (anonymous):

all right so x^2+3xy+y^2=x^2+2xy+y^2=(x+y)^2+xy (x+y)^2 -> x^2+2xy+y^2, and xy what is left

OpenStudy (anonymous):

is it clear now @aimmeBgood ?

OpenStudy (anonymous):

Basically what times what gives you x^2 and what times what gives you 2y^2,